# StreamSplat: Towards Online Dynamic 3D Reconstruction from Uncalibrated Video Streams

### v1

@@ 3.2 Dynamic Deformation Prediction
[…]
where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation. After passing through a few decoder blocks, we obtain the Deformation Embeddings $[\mathbf{d}_{t_{1}},\mathbf{d}_{t_{2}}]\in\mathbb{R}^{2N\times D}$, which are then upsampled via the same upsampler to produce deformation tokens $\hat{\mathbf{d}}\in\mathbb{R}^{32N\times D/4}$. Each deformation token $\hat{\mathbf{d}}_{j}$ is concatenated with the corresponding static Gaussian token from the same frame $\hat{\mathbf{E}}_{j}$ to form a joint token $\hat{\mathbf{E}}_{j}\oplus\hat{\mathbf{d}}_{j}$, which is passed through a 2-layer MLP head to predict the deformation field, including velocity $\mathbf{v}_{j}$ and opacity coefficient $\boldsymbol{\gamma}_{j}$. These deformation fields allow Gaussians to move and fade over time, enabling computation of dynamic 3D Gaussians at arbitrary times for continuous scene reconstruction.
@@ 3.3 Training and Inference
Robust Stage-wise Training. To address the difficulty of jointly optimizing static 3D Gaussian encoding and dynamic deformation prediction, we adopt a two-stage training protocol:
@@
    where $\mathcal{L}_{\text{recon}}$ is the sum of $L_{2}$ loss (RGB space) and LPIPS loss [72], and $\mathcal{L}_{\text{depth}}$ is a scale- and shift-invariant depth loss [73]:
@@
    where $\mathcal{L}_{\text{mask}}$ is an auxiliary reconstruction loss that encourages the model to focus on moving foreground regions using a binary segmentation mask $M_{t}$ from datasets [33, 34].
Online Inference Pipeline. As illustrated in Figure 2, our StreamSplat can instantly process an incoming video stream $(I_{0},I_{1},\dots,I_{N})$ frame-by-frame. Let $\mathbf{h}_{t-1}$ and $\mathbf{f}_{t-1}$ denote the 3DGS Embedding and DINO feature from the previous frame (initialized as zero for $t=0$). When a new frame $I_{t}$ arrives, we estimate its pseudo-depth $D_{t}$ [65] and encode it into 3DGS Embedding $\mathbf{h}_{t}$. These 3DGS Embeddings and DINO features are passed to the dynamic decoder to predict a bidirectional deformation field. The resulting field is applied to both $\mathcal{G}_{t-1}$ and $\mathcal{G}_{t}$, which are fused and rendered from arbitrary viewpoints and interpolated times. Finally, $\mathbf{h}_{t}$ and $\mathbf{f}_{t}$ are cached for the next step.
@@ 4 Experiments
Figure 3: Qualitative results on RE10K. StreamSplat produces detailed and consistent reconstructions across diverse scenes. Videos are provided on the website.
@@ 4.1 Experimental Settings
Training Datasets. We pre-train StreamSplat on a combination of real-world datasets, including static scenes from CO3Dv2 [31] and RealEstate10K (RE10K) [32], and dynamic video datasets DAVIS [33] and YouTube-VOS [34]. CO3Dv2 and RE10K are treated as pure video datasets without using any pre-calibrated camera information. For dynamic scenes, we utilize object segmentation masks from DAVIS and YouTube-VOS to supervise motion-aware components, and no mask supervision is applied to CO3Dv2 and RE10K. We follow the official train/validation splits and train only on the training sets. The same pre-trained model is evaluated across both static and dynamic benchmarks.
Implementation Details. StreamSplat is trained on 8 NVIDIA A100 GPUs for approximately 3 days. We use FlashAttention-2 [74], gradient checkpointing [63], and mixed-precision training with BF16 for better efficiency. Input frames are resized to $512\times 288$ to preserve aspect ratio for pixel-aligned 3DGS prediction. We apply image-level augmentations following EDM [75] and optimize using AdamW [76] with gradient clipping set to 1.0. For Stage 1, we use a batch size of 128, a peak learning rate of $5\times 10^{-4}$ with 20K linear warm-up iterations, and weight decay of 0.05. For Stage 2, we use a batch size of 256 (with gradient accumulation), a peak learning rate of $1\times 10^{-4}$ with 100K linear warm-up iterations, and weight decay of 0.05. Additional training details, including hyperparameters and model configurations, are provided in the Appendix B.
Figure 4: Qualitative comparison on DAVIS. Blue box: given frames; Red box: interpolated frames. Our StreamSplat produces high-fidelity and temporal coherent interpolations across both (a) 5-frame and (b) 8-frame interval tasks. Videos are provided on the website.
Evaluation Settings. We follow prior works [15, 28] and report peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) [77], and LPIPS [72], all evaluated at a resolution of $256\times 256$ for fair comparison [15, 28]. For static scene reconstruction, we follow [15], using two randomly sampled input views with at least $60\%$ overlap and five target views sampled between them. For dynamic scene reconstruction, we first evaluate on full video sequences. To further assess dynamic modeling, we follow [28] and evaluate video interpolation on subsampled sequences with 5-frame and 8-frame intervals and use intermediate frames for metric computation. Our website provide video results across all four datasets, including extended demonstrations on long videos.
Method Rep. Type Given View Novel View Average
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
pixelSplat [15] Stat. 30.70 0.952 0.055 28.31 0.905 0.097 28.99 0.918 0.085
MVSplat [55] 31.48 0.962 0.046 28.48 0.909 0.091 29.34 0.924 0.078
NoPoSplat [16] 29.50 0.939 0.069 28.65 0.913 0.096 28.90 0.920 0.089
CoDeF [35] Dyn. 35.13 0.943 0.091 20.51 0.591 0.402 21.77 0.625 0.374
DGMarbles [78] 27.48 0.867 0.232 23.40 0.727 0.333 23.73 0.738 0.325
@@
Method Scene Rep. Key Frames Middle-4 Frames Time
@@
Omnimotion [26] NeRF 24.11 0.714 0.371 > 8 hrs
RoDynRF [79] NeRF 24.79 0.723 0.394 > 24 hrs
CoDeF [35] NeRF 31.49 0.939 0.088 19.40 0.498 0.400 ~10 mins
MonST3R [14] Points 42.33 0.980 0.012 ~30 s
4DGS [45] 3DGS 18.12 0.573 0.513 ~40 mins
Splater a Video [30] 3DGS 28.63 0.837 0.228 ~30 mins
DGMarbles [78] 3DGS 28.38 0.879 0.172 21.33 0.619 0.313 ~30 mins
StreamSplat (Ours) 3DGS 37.83 0.982 0.016 23.66 0.684 0.193 1.48 s
@@ 4.2 Static Scene Reconstruction
We evaluate StreamSplat on the RE10K benchmark and compare it with recent static and dynamic reconstruction methods. Results on CO3Dv2 are in Appendix C. For dynamic methods without camera pose input, novel views are rendered using relative timestamps. Quantitative and qualitative results are presented in Table 1 and Figures 3 and 7. StreamSplat significantly outperforms all static baselines on the given-view reconstruction task. However, on novel-view reconstruction, static methods still lead, although StreamSplat achieves the best performance among all dynamic approaches.
According to our qualitative comparison in Figures 3 and 7, we attribute this to the absence of accurate camera poses input in our model. Our StreamSplat assumes smooth camera motion between two input views and relies solely on relative timestamps for novel view synthesis, which can cause slight misalignments. Furthermore, unlike static models that assume a fixed scene, StreamSplat is designed to model scene dynamics, which may introduce motion artifacts that hinder performance under static benchmarks. Despite these challenges, our StreamSplat achieves competitive average performance across all test views and consistently surpasses all dynamic baselines in every evaluation setting.
4.3 Dynamic Scene Reconstruction
Figure 5: Visualization of reconstructed dynamic scene from canonical and novel views. Our method captures consistent 3D motion over time, enabling faithful reconstruction at arbitrary time and viewpoints. Videos are provided on the website.
Table 3: Quantitative results on the DAVIS-7 [28] with 8-frame interval.
Method Type PSNR SSIM LPIPS
AMT [80] pixel 21.09 0.544 0.254
RIFE [81] 20.48 0.511 0.258
FILM [82] 20.71 0.528 0.270
LDMVFI [83] 19.98 0.479 0.276
VIDIM [28] 19.62 0.470 0.257
CoDeF [35] 3D 20.34 0.520 0.365
DGMarbles [78] 19.83 0.548 0.353
Ours 22.10 0.613 0.234
We evaluate StreamSplat on the DAVIS benchmark and compare it with state-of-the-art methods, and report the main results in Table 2. Notably, StreamSplat is the only method capable of near real-time dynamic 3D reconstruction per frame.
For key-frame reconstruction task, StreamSplat achieves performance competitive with the state-of-the-art scene-coordinate-based method MonST3R [14], which represents the dynamic scenes as sequences of static 3D point clouds. However, MonST3R requires extensive post-optimization and is limited to key-frame reconstruction. In contrast, StreamSplat operates in near real time and explicitly models the scene dynamics, enabling reconstruction of intermediate frames across substantial temporal gaps.
We further evaluate dynamic modeling performance under 5-frame and 8-frame interval settings, comparing against optimization-based 3D reconstruction methods and pixel-level video-interpolation methods. As shown in Table 3 and Figures 4 and 8, StreamSplat consistently outperforms all baselines, including video-interpolation methods which lack explicit 3D modeling. This highlights the effectiveness of our approach in modeling temporally coherent dynamic scenes. Moreover, as illustrated in Figures 4 and 8, StreamSplat maintains high visual fidelity even under challenging scenarios such as large camera motion and reflective surfaces, where other methods often fail.
In addition, we demonstrate the robustness of StreamSplat in long-range dynamic modeling. As shown in Figure 5, our method maintains coherent 3D structure and appearance across large spatio-temporal distances, from both canonical and novel viewpoints.
@@ 4.4 Ablation Studies
We conduct ablation studies to evaluate the contribution of each key component in StreamSplat. Quantitative and qualitative results are presented in Table 4 and Figures 6 and 9.
Ablation on probabilistic position sampling. We evaluate the impact of probabilistic position sampling for 3DGS by comparing it with a deterministic counterpart. As shown in Table 4, probabilistic sampling yields a substantial improvement (6.36dB) in PSNR for key-frame reconstruction. Qualitatively, the deterministic variant is prone to local minima, particularly along the depth axis, resulting in blurry and inaccurate reconstructions. This aligns with findings from prior work [21, 15] and highlights the importance of probabilistic position prediction in feed-forward 3DGS models.
@@
Ablation on bidirectional deformation field. We evaluate the effectiveness of the bidirectional deformation field by comparing it with a conventional deformation field [44] on middle-frame reconstruction. As shown in Table 4 and Figure 9, the bidirectional variant significantly improves reconstruction quality. The baseline struggles to preserve pixel-aligned structures, leading to noticeable error accumulation over longer sequences.
@@ 5 Conclusion
In this paper, we introduced StreamSplat, the first feed-forward framework for instant, online dynamic 3D reconstruction from uncalibrated video streams. By incorporating a probabilistic position sampling strategy and a bidirectional deformation field with adaptive Gaussian fusion, our StreamSplat effectively addresses key challenges in online dynamic reconstruction, allowing to produce accurate dynamic 3D Gaussian Splatting (3DGS) representations from arbitrarily long video streams. These representations faithfully capture scene dynamics and support interpolation at arbitrary time. Extensive experiments on both static and dynamic benchmarks validate the superior performance of StreamSplat in terms of reconstruction quality and dynamic scene modeling. In future, we plan to explore its potential in video generation and autonomous driving.
Acknowledgement
This work was funded, in part, by the NSERC DG Grant (No. RGPIN-2022-04636, No. RGPIN-2019-05448), the Vector Institute for AI, Canada CIFAR AI Chair, and a Google Gift Fund. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through the Digital Research Alliance of Canada alliance.can.ca, and companies sponsoring the Vector Institute www.vectorinstitute.ai/#partners, and Advanced Research Computing at the University of British Columbia. Additional hardware support was provided by John R. Evans Leaders Fund CFI grant. ZW and QY are supported by UBC Four Year Doctoral Fellowships. We thank Qiuhong Shen for constructive discussions and helpful comments.
@@ Appendix A Background
3D Gaussian Splatting. 3D Gaussian Splatting (3DGS) [21] represents a static scene using a collection of 3D Gaussians. Each Gaussian $\mathcal{G}$ is defined by its mean (position) $\boldsymbol{\mu}$, covariance matrix $\mathbf{\Sigma}$, opacity $\boldsymbol{\alpha}$, and color represented by spherical harmonics (SH) coefficients $\boldsymbol{c}$. The final opacity of a 3D Gaussian at a given point $\mathbf{x}$ is computed as:
$\boldsymbol{\alpha}(\mathbf{x})=\boldsymbol{\alpha}\cdot\exp\left(-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x}-\boldsymbol{\mu})\right).$ (6)
@@
The 3D Gaussians are projected using the approximate transformation [29]:
@@
The pixel color $\boldsymbol{C}$ is obtained by alpha-blending the projected 2D Gaussians sorted by depth:
$\boldsymbol{C}=\sum_{i=1}^{N}\boldsymbol{c}_{i}\boldsymbol{\alpha}_{i}\prod_{j=1}^{i-1}(1-\boldsymbol{\alpha}_{j})$ (11)
where $\boldsymbol{\alpha}_{i}$ is the opacity of the $i$-th projected Gaussian obtained by Eq. (6).
Dynamic 3D Gaussian Splatting. Dynamic 3D Gaussian Splatting extends the original 3DGS with a deformation field to model the motion of the Gaussians over time [43]. Typically, the deformation fields $\mathcal{D}(t)=\{\boldsymbol{\mu}(t),\mathbf{R}(t),\boldsymbol{\alpha}(t)\}$ are used to update the static canonical Gaussian with various of approaches [44, 45, 30, 47, 48]. For example, given a static position $\mu$ and a deformation field $\mu(t)$, the deformed position can be obtained as: $\hat{\mu}_{t}=\mu+\mu(t)$.
Figure 7: Qualitative results on RE10K. Blue box: given frames; Red box: interpolated frames. Compared to other dynamic-scene reconstruction methods, StreamSplat produces more detailed and consistent 3D reconstructions across diverse scenes, whereas other methods often exhibit distortions in both color and geometry.
Input : Gaussian token $\hat{\mathbf{E}}$, deformation token $\hat{\mathbf{d}}$, current time $t$, frame time $t_{0}$
Output : 3DGS Parameters: position $\boldsymbol{\mu}$, scale $\mathbf{S}$, rotation $\mathbf{R}$, opacity $\boldsymbol{\alpha}$, color $\boldsymbol{c}$
// Static Prediction (Time-Invariant Parameters)
$\mathbf{f}_{\text{static}}\leftarrow\textsc{Linear}(\hat{\mathbf{E}})$ ;
  // Static feature
$(\mathbf{f}_{\mu_{p}},\;\mathbf{f}_{\Sigma_{p}},\;\mathbf{f}_{S},\;\mathbf{f}_{R},\;\mathbf{f}_{\alpha},\;\mathbf{f}_{c})\leftarrow\textsc{Split}(\mathbf{f}_{\text{static}})$ ;
$\boldsymbol{o}\sim\mathcal{N}_{[-1,1]}(\tanh(\mathbf{f}_{\mu_{p}}),\exp^{2}(\mathbf{f}_{\Sigma_{p}}))$ ;
  // Sample position offset
$\boldsymbol{\mu}_{0}\leftarrow\left[u+\boldsymbol{o}_{0},\;v+\boldsymbol{o}_{1},\;\dfrac{1}{0.5+0.5\cdot\boldsymbol{o}_{2}}\right]$ ;
  // Pixel-aligned position
$\boldsymbol{\alpha}_{0}\leftarrow\textsc{Sigmoid}(\mathbf{f}_{\alpha})$ ;
$\mathbf{S}\leftarrow 0.1*\textsc{Softplus}(\mathbf{f}_{S})$ ;
$\mathbf{R}\leftarrow\textsc{Normalize}(\mathbf{f}_{R}$) ;
$\boldsymbol{c}\leftarrow\textsc{Sigmoid}(\mathbf{f}_{c})$ ;
// Deformation Prediction (Time-Dependent Parameters)
$\mathbf{f}_{\text{def}}\leftarrow\textsc{MLP}\left(\textsc{Linear}(\hat{\mathbf{E}})\oplus\hat{\mathbf{d}}\right)$ ;
  // Deformation feature
$(\Delta\mathbf{f}_{\mu_{p}},\;\Delta\mathbf{f}_{\Sigma_{p}},\;\Delta\mathbf{f}_{\gamma})\leftarrow\textsc{Split}(\mathbf{f}_{\text{def}})$ ;
$\mathbf{v}\sim\mathcal{N}_{[-1,1]}(\tanh(\Delta\mathbf{f}_{\mu_{p}}),\exp^{2}(\Delta\mathbf{f}_{\Sigma_{p}}))$ ;
  // Sample position velocity
$\boldsymbol{\gamma}\leftarrow\left[\textsc{ReLU}(\Delta\mathbf{f}_{\gamma_{0}}),\textsc{Sigmoid}(\Delta\mathbf{f}_{\gamma_{1}})\right]$ ;
// Apply time-dependent translation and opacity decay
$\boldsymbol{\mu}\leftarrow\boldsymbol{\mu}_{0}+\mathbf{v}\cdot t$ ;
$\boldsymbol{\alpha}\leftarrow\boldsymbol{\alpha}_{0}\cdot\dfrac{\textsc{Sigmoid}\left(-\boldsymbol{\gamma}_{0}\left(|t-t_{0}|-\boldsymbol{\gamma}_{1}\right)\right)}{\textsc{Sigmoid}\left(\boldsymbol{\gamma}_{0}\cdot\boldsymbol{\gamma}_{1}\right)}$ ;
  // Eq. (1)
return $\boldsymbol{\mu},\mathbf{S},\mathbf{R},\boldsymbol{\alpha},\boldsymbol{c}$
Algorithm 1 3DGS Parameters Predictions

### v2

@@ 3.2 Dynamic Deformation Prediction
2pt return $\tilde{\mathcal{G}}(t)$, $\mathbf{h}_{k}$, $\hat{I}_{t_{k-1}\rightarrow t_{k}}$
Algorithm 2 Online Inference
@@ 3.3 Training and Inference
Robust Stage-wise Training. To address the difficulty of jointly optimizing static 3D Gaussian encoding and dynamic deformation prediction, we adopt a two-stage training protocol.
@@
where $\mathcal{L}_{\text{recon}}$ is the sum of $L_{2}$ loss (RGB space) and LPIPS loss Zhang et al. (2018), and $\mathcal{L}_{\text{depth}}$ is a scale- and shift-invariant depth loss Ranftl et al. (2020):
@@
where $\mathcal{L}_{\text{mask}}$ is an auxiliary reconstruction loss that encourages the model to focus on moving foreground regions using a binary segmentation mask $M_{t}$ from datasets Pont-Tuset et al. (2017); Xu et al. (2018).
Online Inference Pipeline. Algorithm 2 summarizes the online inference. We maintain a canonical Gaussian set $\tilde{\mathcal{G}}(t)$ and store the previous embedding $\mathbf{h}_{k-1}$. For each incoming frame $I_{k}$, we predict pseudo-depth $D_{k}$, encode $\mathbf{h}_{k}$, and use the Dynamic Decoder with Algorithm 1 on $(\mathbf{h}_{k-1},\mathbf{h}_{k})$ to produce two Gaussian sets: $\mathcal{G}_{k-1}^{+}(t)$ (previous Gaussians deformed forward) and $\mathcal{G}_{k}^{-}(t)$ (current Gaussians deformed backward). We then (i) Update $\tilde{\mathcal{G}}(t)$ with $\mathcal{G}_{k-1}^{+}(t)$, setting the active deformation of matched 3DGS primitives to the new forward field, (ii) aggregate and fuse $\mathcal{G}_{k}^{-}(t)$, (iii) render frames $\hat{I}_{t_{k-1}\!\rightarrow t_{k}}$ from given viewpoints, and (iv) prune Gaussians whose opacity at $t_{k}$ decays to zero. In the end, we store $\mathbf{h}_{k}$ for the next step.
Figure 4: Persistent Gaussians across frames. Red/green-marked Gaussians from initial frame are propagated across frames, showing that adaptive Gaussian fusion preserves long-term temporal consistency under viewpoint and motion changes. Videos are available on the project website.
@@ 4.1 Experimental Settings
Training Datasets. We pre-train StreamSplat on a combination of real-world datasets, including static scenes from CO3Dv2 Reizenstein et al. (2021) and RealEstate10K (RE10K) Zhou et al. (2018), and dynamic video datasets DAVIS Pont-Tuset et al. (2017) and YouTube-VOS Xu et al. (2018). CO3Dv2 and RE10K are treated as pure video datasets without using any pre-calibrated camera information. For dynamic scenes, we utilize object segmentation masks from DAVIS and YouTube-VOS to supervise motion-aware components, and no mask supervision is applied to CO3Dv2 and RE10K. We follow the official train/validation splits and train only on the training sets. The same pre-trained model is evaluated across both static and dynamic benchmarks.
Implementation Details. StreamSplat is trained on 8 NVIDIA A100 GPUs for approximately 3 days. We use FlashAttention-2 Dao (2023), gradient checkpointing Chen et al. (2016), and mixed-precision training with BF16 for better efficiency. Input frames are resized to $512\times 288$ to preserve aspect ratio for pixel-aligned 3DGS prediction. We apply image-level augmentations following EDM Karras et al. (2022) and optimize using AdamW Loshchilov and Hutter (2017) with gradient clipping set to 1.0. For Stage 1, we use a batch size of 128, a peak learning rate of $5\times 10^{-4}$ with 20K linear warm-up iterations, and weight decay of 0.05. For Stage 2, we use a batch size of 256 (with gradient accumulation), a peak learning rate of $1\times 10^{-4}$ with 100K linear warm-up iterations, and weight decay of 0.05. Additional training details are in Appendix B.
Figure 5: Qualitative comparison on DAVIS. Blue box: given frames; Red box: novel frames. StreamSplat produces high-fidelity and temporal coherent results across both (a) 5-frame and (b) 8-frame interval tasks.
Evaluation Settings. We follow prior works Charatan et al. (2024); Jain et al. (2024) and report peak signal-to-noise ratio (PSNR), structural similarity index (SSIM) Wang et al. (2004), and LPIPS Zhang et al. (2018), all evaluated at a resolution of $256\times 256$ for fair comparison Charatan et al. (2024); Jain et al. (2024). For static scene reconstruction, we follow Charatan et al. (2024), using two randomly sampled input views with at least $60\%$ overlap and five target views sampled between them. For dynamic scene reconstruction, we evaluate on sparse, uncalibrated videos subsampled at 5- and 8-frame intervals Jain et al. (2024) and compute metrics on all non-input frames. We also conducted a zero-shot evaluation on DyCheck Gao et al. (2022) and NVIDIA Dynamic Scene Yoon et al. (2020) in Appendix C. Videos are available on the project website.
@@
Method Scene Rep. Key Frames Middle‑4 Frames Time
@@
Omnimotion Wang et al. (2023) NeRF 24.11 0.714 0.371 > 8 hrs
RoDynRF Liu et al. (2023) NeRF 24.79 0.723 0.394 > 24 hrs
CoDeF Ouyang et al. (2024) NeRF 31.49 0.939 0.088 19.40 0.498 0.400 ~10 mins
MonST3R Zhang et al. (2024a) Points 42.33 0.980 0.012 ~30 s
4DGS Duan et al. (2024) 3DGS 18.12 0.573 0.513 ~40 mins
Splater a Video Sun et al. (2024) 3DGS 28.63 0.837 0.228 ~30 mins
DGMarbles Stearns et al. (2024) 3DGS 28.38 0.879 0.172 21.33 0.619 0.313 ~30 mins
StreamSplat (Ours) 3DGS 37.83 0.982 0.016 23.66 0.684 0.193 ~1.48 s
Figure 6: Visualization of reconstructed scene from canonical and novel views. Our method captures consistent 3D scene over time, enabling faithful reconstruction at arbitrary time and viewpoints.
Figure 7: Qualitative results on RE10K. Blue box: given frames; Red box: novel frames. StreamSplat produces consistent reconstructions, while others exhibit color/geometry distortions.
4.2 Dynamic Scene Reconstruction
Table 2: 8-frame interval interpolation results.
Method Type PSNR SSIM LPIPS
AMT 2D 21.09 0.544 0.254
RIFE 20.48 0.511 0.258
FILM 20.71 0.528 0.270
LDMVFI 19.98 0.479 0.276
VIDIM 19.62 0.470 0.257
CoDeF 4D 20.34 0.520 0.365
DGMarbles 19.83 0.548 0.353
Ours 22.10 0.613 0.234
Video Reconstruction. We evaluate StreamSplat on the DAVIS benchmark and compare it with state-of-the-art methods, and report the main results in Table 1. Notably, StreamSplat is the only method capable of near real-time dynamic 3D reconstruction, with a runtime of only 0.049s per-frame and a 1200× speedup over optimization-based baselines.
For key-frame reconstruction task, StreamSplat achieves performance competitive with the state-of-the-art scene-coordinate-based method MonST3R Zhang et al. (2024a), which represents the dynamic scenes as sequences of static 3D point clouds. However, MonST3R requires extensive post-optimization and is limited to key-frame reconstruction. In contrast, StreamSplat operates in near real time and explicitly models the scene dynamics, enabling reconstruction of intermediate frames across substantial temporal gaps.
We further evaluate dynamic modeling performance under 5-frame and 8-frame interval settings, comparing against optimization-based 3D reconstruction methods (CoDeF Ouyang et al. (2024), DGMarbles Stearns et al. (2024)) and 2D video-interpolation methods (AMT Li et al. (2023), RIFE Huang et al. (2022), FILM Reda et al. (2022), VIDIM Jain et al. (2024)). As shown in Table 2, Figures 5 and 12, StreamSplat consistently outperforms all baselines, including video-interpolation methods which lack explicit 3D modeling. This highlights the effectiveness of our approach in modeling temporally coherent dynamic scenes. Moreover, as illustrated in Figures 5 and 12, StreamSplat maintains high visual fidelity even under challenging scenarios such as large camera motion and reflective surfaces, where other methods often fail.
In addition, we demonstrate the robustness of StreamSplat in long-range dynamic modeling and novel view synthesis. As shown in Figure 6, our method maintains coherent 3D structure and appearance across large spatio-temporal distances, from both canonical and novel viewpoints.
Temporal Coherence and Persistency. We evaluate the temporal coherence of our modeled representation by color-coding random Gaussians in the initial frame and propagating their identities through the fused Gaussians over time (Figures 4 and 15). Despite viewpoint/scale changes, motion, occlusions, and blur, the highlighted Gaussians remain stable, indicating that our adaptive fusion maintains long-term consistency without explicit matching or iterative fusion.
Table 3: RE10K results. PSNR/LPIPS for 2 given views and 5 novel views.
Method Given View Novel View Average
CoDeF 35.13 / 0.09 20.51 / 0.40 21.77 / 0.37
DGMarbles 27.48 / 0.23 23.40 / 0.33 23.73 / 0.32
StreamSplat 41.60 / 0.01 24.68 /0.167 29.51 / 0.12
@@ 4.3 Static Scene Reconstruction
We evaluate StreamSplat on RE10K against recent dynamic reconstruction methods; results for static methods and CO3Dv2 are deferred to Appendix C. For pose-free dynamic pipelines, novel views are rendered using relative timestamps. Quantitative and qualitative results are presented in Table 3 and 5, Figures 7, 11 and 10. StreamSplat significantly outperforms all static baselines on the given-view reconstruction, and consistently outperforms all dynamic baselines in every evaluation setting, with an average gain of 5.78dB in PSNR. Qualitatively, in Figure 7, StreamSplat produces more detailed and consistent 3D reconstructions across diverse scenes, whereas other dynamic reconstruction methods often exhibit distortions in both color and geometry.
@@ 4.4 Ablation Studies
We conduct ablation studies to evaluate the contribution of each key component in StreamSplat. Quantitative and qualitative results are presented in Table 4 and Figures 8 and 13.
Ablation on probabilistic position sampling. We evaluate the impact of probabilistic position sampling for 3DGS by comparing it with a deterministic counterpart. As shown in Table 4, probabilistic sampling yields a substantial improvement (6.36dB) in PSNR for key-frame reconstruction. Qualitatively, the deterministic variant is prone to local minima, particularly along the depth axis, resulting in blurry and inaccurate reconstructions. This aligns with findings from prior work Kerbl et al. (2023); Charatan et al. (2024) and highlights the importance of probabilistic position prediction in feed-forward 3DGS models.
@@
Ablation on bidirectional deformation field. We evaluate the effectiveness of the bidirectional deformation field by comparing it with a conventional deformation field Li et al. (2024) on middle-frame reconstruction. As shown in Table 4 and Figure 13, the bidirectional variant significantly improves reconstruction quality. The baseline struggles to preserve pixel-aligned structures, leading to noticeable error accumulation over longer sequences.
@@ 5 Conclusion
In this paper, we introduced StreamSplat, a fully feed-forward framework for instant, online dynamic 3D reconstruction from uncalibrated video streams. By incorporating a probabilistic position sampling strategy and a bidirectional deformation field with adaptive Gaussian fusion, our StreamSplat effectively addresses key challenges in online dynamic reconstruction, allowing to produce accurate dynamic 3D Gaussian Splatting (3DGS) representations from arbitrarily long video streams. These representations faithfully capture scene dynamics and support rendering at arbitrary time and viewpoints. Extensive experiments on both static and dynamic benchmarks validate the superior performance of StreamSplat in terms of reconstruction quality and dynamic scene modeling. In future, we plan to explore its potential in video generation and autonomous driving.
Acknowledgements
This work was supported, in part, by the NSERC DG Grant (No. RGPIN-2022-04636, No. RGPIN-2019-05448), the Vector Institute for AI, the Canada CIFAR AI Chair program, a Google Gift Fund, and a NVIDIA Academic Grant. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through the Digital Research Alliance of Canada www.alliancecan.ca, and companies sponsoring the Vector Institute www.vectorinstitute.ai/#partners, and Advanced Research Computing at the University of British Columbia. Additional resource was provided by John R. Evans Leaders Fund CFI grant. ZW and QY are supported by UBC Four Year Doctoral Fellowships. We thank Qiuhong Shen and Qingshan Xu for constructive discussions and helpful comments.
@@ Appendix A Background
3D Gaussian Splatting. 3D Gaussian Splatting (3DGS) Kerbl et al. (2023) represents a static scene using a collection of 3D Gaussians. Each Gaussian $\mathcal{G}$ is defined by its mean (position) $\bm{\mu}$, covariance matrix $\mathbf{\Sigma}$, opacity $\bm{\alpha}$, and color represented by spherical harmonics (SH) coefficients $\bm{c}$. The final opacity of a 3D Gaussian at a given point $\mathbf{x}$ is computed as:
$\bm{\alpha}(\mathbf{x})=\bm{\alpha}\cdot\exp\left(-\frac{1}{2}(\mathbf{x}-\bm{\mu})^{T}\mathbf{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})\right).$ (5)
@@
The 3D Gaussians are projected using the approximate transformation Zwicker et al. (2001):
@@
The pixel color $\bm{C}$ is obtained by alpha-blending the projected 2D Gaussians sorted by depth:
$\bm{C}=\sum_{i=1}^{N}\bm{c}_{i}\bm{\alpha}_{i}\prod_{j=1}^{i-1}(1-\bm{\alpha}_{j})$ (10)
where $\bm{\alpha}_{i}$ is the opacity of the $i$-th projected Gaussian obtained by Eq. 5.
Dynamic 3D Gaussian Splatting. Dynamic 3D Gaussian Splatting extends the original 3DGS with a deformation field to model the motion of the Gaussians over time Luiten et al. (2024). Typically, the deformation fields $\mathcal{D}(t)=\{\bm{\mu}(t),\mathbf{R}(t),\bm{\alpha}(t)\}$ are used to update the static canonical Gaussian with various of approaches Li et al. (2024); Duan et al. (2024); Sun et al. (2024); Lee et al. (2024); Shi et al. (2025). For example, given a static position $\mu$ and a deformation field $\mu(t)$, the deformed position can be obtained as: $\hat{\mu}_{t}=\mu+\mu(t)$.
Figure 9: StreamSplat framework. Given a pair of frames, we first encode them using the Static Encoder to produce canonical 3D Gaussians (Section 3.1), and then pass the 3DGS Embeddings to the Dynamic Decoder to predict the deformation field (Section 3.2). The resulting dynamic 3D Gaussians can be rendered at arbitrary time to produce RGB images and depth maps.